OpenStudy (anonymous):

PLEASE help me, I need help multiplying rational expressions: 9s/-1 * 1/-9s

OpenStudy (anonymous):

\[\frac{ 9s }{ -1 } * \frac{ 1 }{ -9s }\]

OpenStudy (immanuelv):

\[9s * \frac{ 1 }{ -9s}\] \[s = 1\] everywhere except 0 \[(-\infty,0) \cup (0, \infty)\]

OpenStudy (anonymous):

Okay that makes sense, could you help me with another @immanuelv ?

OpenStudy (immanuelv):

sure

OpenStudy (anonymous):

\[\frac{ -2x }{ 3y } * \frac{ -9x^2 }{ 2 }\]

OpenStudy (immanuelv):

\[\frac{ 3x ^{3} }{ y }\] it would be all Reals and y cannot be 0

OpenStudy (anonymous):

Okay I wrote it out, that one makes sense too.

OpenStudy (immanuelv):

im a bit shaky on the domain.... >.< sorry

OpenStudy (immanuelv):

haha glad it worked out for ya

OpenStudy (anonymous):

Thanks!